# Acceleration of multi-component multiple-precision arithmetic with branch-free algorithms and SIMD vectorization

### 16 Mar 2026 v1

@@ 3.2 Benchmark Tests for Complex Matrix Multiplication
Owing to the superiority of BF arithmetic for TD and QD computations demonstrated previously, the complex matrix multiplication benchmarks employed BF-based implementations for comparison. The complex square matrix multiplication is defined using a uniform random variable $r_{u}\in[0,1]$ and a standard normal random variable $r_{n}\in\mathbb{R}$ as
$\begin{split}C&:=AB,\\ A&=[a_{ij}]_{i,j=1,2,...,n},B=[b_{ij}]_{i,j=1,2,...,n}\in\mathbb{R}^{n\times n}\\ a_{ij}&:=\exp(r_{n})(r_{u}-1/2)+\exp(r_{n})(r_{u}-1/2)\sqrt{-1}\\ b_{ij}&:=\exp(r_{n})(r_{u}-1/2)+\exp(r_{n})(r_{u}-1/2)\sqrt{-1}\end{split}$ (2)
Our complex matrix multiplication is implemented using three real matrix multiplications via the 3M method, a variant of the Karatsuba algorithm [10]. Therefore, both the Strassen and block-based algorithms perform three real matrix multiplications. The numerical accuracy, measured as the minimum of the real and imaginary parts, falls within the ranges of 23.2–24.4 significant decimal digits for DD, 39.2–40.8 for TD, and 55.8–56.5 for QD precision.
@@
Figure 2: Computation time (s) of Strassen matrix multiplication: Snapdragon (left), EPYC (right)
The results demonstrate performance characteristics consistent with those observed for real matrix multiplication. In particular, extracting the computation times for $n=2049$ reveals that our implementation achieved speedups significantly exceeding those for real matrix multiplication relative to MPBLAS, as listed in Table 3.
@@ 4 Benchmark Tests for Polynomial Evaluation and Algebraic Equation Solvers
As demonstrated in the matrix multiplication benchmarks, BF arithmetic offers no speedup for DD precision when the conventional sloppy addition and multiplication are employed. In this section, we present benchmark results for polynomial evaluation and an algebraic equation solver based on the Durand–Kerner (DK) method, both of which utilize TD and QD precision arithmetic, for which performance improvements are anticipated.
@@ 4.1 Evaluation of Real-Coefficient Polynomial Functions
For polynomial evaluation, coefficients of the degree-$n$ real-coefficient polynomial $p_{n}(x)=\sum^{n}_{i=0}a_{i}x^{i}$ are generated randomly, and the results are compared among the standard Horner’s method, Estrin’s method, and SIMD-vectorized Estrin’s method.
@@
Figure 3 shows the computation time (in microseconds) for evaluating the real-coefficient polynomial $p_{n}(x)$ at real arguments. The upper panel corresponds to the Snapdragon platform and the lower panel to the EPYC platform; the standard arithmetic (NoBF) is compared against the branch-reduction BF arithmetic (BF). As the degree (or equivalently, the evaluation count corresponding to the problem size) increases, computation time grows, and the effects of branch elimination and vectorization become more pronounced, particularly for high-precision variants such as TD and QD, where operation count and cost are dominant. Across both architectures, a consistent trend is observed wherein BF reduces computation time relative to NoBF for TD and QD precisions.
@@
Figure 4 shows the computation time (in microseconds) for evaluating the real-coefficient polynomial $p_{n}(x)$ at complex arguments. As each complex multiplication and addition decomposes into multiple real operations, computation times tend to be higher than those for real-argument evaluation shown in Fig. 3. By contrast, as the number of operations increases, the impact of conditional branching and normalization steps also grows; therefore, BF arithmetic should in principle remain effective for complex-argument evaluation. Although the Snapdragon results largely conform to this expectation, the EPYC results show slight discernible difference; in particular, the performance difference for TD arithmetic appears to be minimal, even when one exists.
@@ 4.2 Algebraic Equation Solver Based on the DK Method
The method under comparison is the quadratically convergent DK method, a classical simultaneous iterative solver. Denoting the approximate roots after $k$ iterations as
@@
the recurrence relation, expressed using the monic polynomial $q_{n}(x)=p_{n}(x)/a_{n}$, is given by
@@
The Aberth initial values are expressed by
@@
The test problem was the Chebyshev integration problem, in which the polynomial coefficients, with $a_{n}=1$, are computed as follows:
@@
Table 4 presents the benchmark results obtained by applying the DK method with DD, TD, and QD precision arithmetic to the Chebyshev integration problem described previously.
Table 4: Benchmark test for the Chebyshev integration problem ($n=64$)
@@ 5 Conclusion and Future Work
In this study, we applied BF algorithms, which suppress conditional branching, to multi-component multiple-precision arithmetic (DD/TD/QD) and evaluated their performance on CPU SIMD (AVX2 on EPYC) and Arm Neon. Benchmark tests were conducted for real and complex square matrix multiplication, real- and complex-valued evaluation of real-coefficient polynomials, and the Chebyshev integration problem. The results confirm that, for high precision (particularly TD and QD), the combined use of BF conversion and SIMD vectorization is effective and reduces computation time compared with the standard implementation (NoBF). In addition, the results validate that BF conversion offers no benefit for DD precision arithmetic. These findings indicate that achieving performance improvements in multi-component arithmetic requires (i) branch elimination (BF conversion) and (ii) exploitation of instruction-level parallelism (SIMD vectorization) to be combined, with the implementation strategy selected according to the target problem, precision level, and architecture.
The following directions are identified for future work.
@@
    Adaptive implementation selection (autotuning): As the optimal execution path (NoBF/BF, SIMD width, and loop partitioning) varies with the precision level (DD/TD/QD) and problem size, we will introduce a mechanism for selecting the optimal implementation either at runtime or at build time.
@@
    Extension to application problems: The scope of application will be expanded to include high-precision linear algebra problems (QR decomposition, eigenvalue problems, and iterative methods), special function evaluation, numerical integration, and optimization, and the effectiveness in practical settings (convergence behavior, total runtime, and energy efficiency) will be evaluated.
Through these efforts, we aim to realize the accelerated multi-component high-precision arithmetic in a more general and portable form, establishing a high-precision numerical computing infrastructure applicable to a wide range of computational environments from mobile devices to supercomputers.
@@ Acknowledgment
This research was supported by JSPS KAKENHI Grant Number 23K11127.

### 07 May 2026 v2

@@ 3.2 Benchmark Tests for Complex Matrix Multiplication
Owing to the previously demonstrated superiority of BF arithmetic for TD and QD computations, complex matrix multiplication benchmarks employed BF-based implementations for comparison. Complex square matrix multiplication is defined using a uniform random variable $r_{u}\in[0,1]$ and a standard normal random variable $r_{n}\in\mathbb{R}$ as
$\begin{split}C&:=AB,\\ A&=[a_{ij}]_{i,j=1,2,...,n},B=[b_{ij}]_{i,j=1,2,...,n}\in\mathbb{C}^{n\times n}\\ a_{ij}&:=\exp(r_{n})(r_{u}-1/2)+\exp(r_{n})(r_{u}-1/2)\sqrt{-1}\\ b_{ij}&:=\exp(r_{n})(r_{u}-1/2)+\exp(r_{n})(r_{u}-1/2)\sqrt{-1}\end{split}$ (2)
Our complex matrix multiplication was implemented using three real matrix multiplications via the 3M method, a variant of the Karatsuba algorithm [10] as follows:
$\begin{split}C_{1}&:=\mbox{Re}(A)\mbox{Re}(B),\ C_{2}:=\mbox{Im}(A)\mbox{Im}(B)\\ \mbox{Re}(C)&:=C_{1}-C_{2}\\ \mbox{Im}(C)&:=(\mbox{Re}(A)+\mbox{Im}(A))(\mbox{Re}(B)+\mbox{Im}(B))-C_{1}-C_{2}.\end{split}$ (3)
Therefore, both the Strassen- and block-based algorithms perform three real matrix multiplications. The numerical accuracy, measured as the minimum of the real and imaginary parts, falls within the ranges of 23.2–24.4 significant decimal digits for DD, 39.2–40.8 for TD, and 55.8–56.5 for QD precision.
@@
Figure 2: Computation time (s) of complex-valued Strassen matrix multiplication: Snapdragon (left), EPYC (right)
The results demonstrate performance characteristics consistent with those observed for real matrix multiplication. In particular, extracting the computation times for $n=2049$ reveals that our implementation achieved speedups that significantly exceeded those for real matrix multiplication relative to MPBLAS, as listed in Table 5.
@@ 4 Benchmark Tests for Polynomial Evaluation and Algebraic Equation Solvers
As demonstrated by the matrix multiplication benchmarks, BF arithmetic offers no speedup in DD precision when conventional sloppy addition and multiplication are employed. In this section, we present the benchmark results for a polynomial evaluation and an algebraic equation solver based on the Durand–Kerner (DK) method, both of which utilize TD and QD precision arithmetic, for which performance improvements are anticipated. As described previously, we did not use the BF algorithm for the DD evaluation of a polynomial, but used the BF polynomial evaluation for TD and QD arithmetic.
@@ 4.1 Evaluation of Real-Coefficient Polynomial Functions
For the polynomial evaluation, the coefficients of the degree-$n$ real-coefficient polynomial $p_{n}(x)=\sum^{n}_{i=0}a_{i}x^{i}$ were generated randomly, and the results were compared among the standard Horner’s method, Estrin’s method, and SIMD-vectorized Estrin’s method.
@@
Figure 3 shows the computation time (in microseconds) required to evaluate the real-coefficient polynomial $p_{n}(x)$ in real arguments. The upper panel corresponds to the Snapdragon platform and the lower panel to the EPYC platform; the standard arithmetic (NoBF) is compared with branch-reduction BF arithmetic (BF). As the degree (or equivalently, the evaluation count corresponding to the problem size) increases, the computation time increases, and the effects of branch elimination and vectorization become more pronounced, particularly for high-precision variants such as TD and QD, where the operation count and cost are dominant. Across both architectures, a consistent trend is observed: BF reduces the computation time relative to NoBF for TD and QD precisions.
@@
Figure 4 shows the computation time (in microseconds) required to evaluate the real-coefficient polynomial $p_{n}(x)$ for the complex arguments. Because each complex multiplication and addition decomposes into multiple real operations, the computation times are higher than those for real-argument evaluation, as shown in Fig.  3. In contrast, as the number of operations increases, the impact of the conditional branching and normalization steps also increases; therefore, BF arithmetic should, in principle, remain effective for complex-argument evaluation. Although the Snapdragon results largely conform to this expectation, the EPYC results show a slight discernible difference. In particular, the performance difference for TD arithmetic appears to be minimal, even when one exists.
@@ 4.2 Algebraic Equation Solver Based on the DK Method
The method being compared is the quadratically convergent DK method, which is a classical simultaneous iterative solver. The approximate roots after $k$ iterations are expressed as follows:
@@
The recurrence relationship, represented using the monic polynomial $q_{n}(x)=p_{n}(x)/a_{n}$, is expressed by
@@
The Aberth initial values are expressed as
@@
The test problem was the Chebyshev integration problem, in which the polynomial coefficients with $a_{n}=1$ were computed as follows:
@@
Table 6 presents the benchmark results obtained by applying the DK method with DD, TD, and QD precision arithmetic to the Chebyshev integration problem, as described previously.
Table 6: Benchmark test (unit: second) for the Chebyshev integration problem ($n=64$)
@@ 5 Conclusion and Future Work
In this study, we applied BF algorithms that suppress conditional branching to multicomponent multiple-precision arithmetic (DD/TD/QD) and evaluated their performance on CPU SIMD (AVX2 on EPYC) and Arm Neon. Benchmark tests were conducted for real and complex square matrix multiplications, real- and complex-valued evaluations of real-coefficient polynomials, and the Chebyshev integration problem. The results confirm that, for high precision (particularly TD and QD), the combined use of BF conversion and SIMD vectorization is effective and reduces the computation time compared with the standard implementation (NoBF). In addition, the results validate that BF conversion offers no benefit to DD precision arithmetic. These findings indicate that achieving performance improvements in multicomponent arithmetic requires (i) branch elimination (BF conversion) and (ii) the exploitation of instruction-level parallelism (SIMD vectorization) to be combined with an implementation strategy selected according to the target problem, precision level, and architecture.
The following directions are identified for future research:
@@
    Adaptive implementation selection (autotuning): As the optimal execution path (NoBF/BF, SIMD width, and loop partitioning) varies with the precision level (DD/TD/QD) and problem size, we will introduce a mechanism for selecting the optimal implementation at either runtime or build time.
@@
    Extension to application problems: The scope of the application will be expanded to include high-precision linear algebra problems (QR decomposition, eigenvalue problems, and iterative methods), special function evaluation, numerical integration, and optimization, and the effectiveness in practical settings (convergence behavior, total runtime, and energy efficiency) will be evaluated.
Through these efforts, we aim to realize accelerated multicomponent high-precision arithmetic in a more general and portable form by establishing a high-precision numerical computing infrastructure applicable to a wide range of computational environments, ranging from mobile devices to supercomputers.
@@ Acknowledgment
This research was supported by JSPS KAKENHI (grant number 23K11127).